# Shuttling Compiler for Trapped-Ion Quantum Computers Based on Large Language Models

### v1

@@ 4.2.2 Unseen architectures
In addition to linear and branched one-dimensional architectures with junctions, we evaluated the generalization capabilities of the fine-tuned LLMs on the unseen architectures shown in Fig. 5. While the layouts in 5(a) and 5(b) feature three-way junctions, similar to the branched architectures used during training, the architecture in 5(c) includes four-way junctions.
For the architectures shown in 5(a) and 5(b), none of the models generated a valid shuttling schedule for any of the evaluated circuits. While on the first architecture, all models failed immediately during schedule generation, on the second one, the models were able to execute up to four gates before failing.
In contrast, one model succeeded in generating a valid schedule for the architecture shown in 5(c) for the circuit 4mod5-bdd_287. LLaMA at $T=0.7$ produced a valid schedule with 569 shuttling operations but required 55 retries, increasing the total token count by a factor of 6.0 relative to the final schedule. At $T=1.0$, the model executed 74 of the 106 gates before failing. All other models failed to generate valid schedules on this architecture. Nevertheless, partial progress was observed: Qwen-14B executed up to 52 gates, Qwen-4B up to 33 gates, and Gemma at $T=1.0$ 16 gates, while Gemma at $T=0.7$ and DeepSeek failed at an early stage.
Consistent with the results on linear and branched architectures, performance on the four-way junction architecture degraded as the number of qubits increased. For the circuit mini_alu_305, no model generated a valid schedule. LLaMA executed up to 48 gates before failing, whereas both Qwen models and Gemma executed approximately 15 gates. DeepSeek performed worst, failing to execute even a single gate. Excluding DeepSeek, these results indicate that limited progress remains possible for 10-qubit circuits, although complete schedules cannot be reliably generated.
@@
Overall, our results show that fine-tuned LLMs can generate valid shuttling schedules and can, in some cases, outperform classical compilers. However, producing complete schedules becomes increasingly challenging as the number of qubits and gates increases. The models frequently enter invalid states, requiring retries that substantially inflate token counts without contributing to the final output. Generalization to previously unseen architectures remains limited. Across all experiments, LLaMA demonstrated the most stable performance, whereas Qwen, DeepSeek, and Gemma failed more frequently and at earlier stages as circuit size increased. When comparing the both Qwen models, it can be noticed that the smaller model with 4 billion parameters produced better results than the the larger model with 14 billion parameters. This indicates, together with LLaMA which is with 3 billion parameters the smallest model we tested, that smaller models can be better adapted to shuttling schedule generation than larger ones.
@@ 5 Conclusion and Outlook
We have presented a shuttling compiler for trapped-ion quantum computers based on large language models (LLMs). We fine-tuned several LLMs on datasets derived from shuttling schedules generated by classical compilers for linear and branched one-dimensional trap architectures and evaluated them on circuits of varying sizes. For small circuits on both linear and junction layouts, the models were able to generate valid schedules and in some cases even required fewer operations than the classical compilers.
As circuit size increased, generating complete shuttling schedules became substantially more difficult. Larger numbers of qubits make trap reconfiguration increasingly complex and often require long, recursive sequences of shuttling operations. As a result, the models frequently succeeded in generating valid operations only for an initial portion of the circuit before reaching trap configurations from which continuation would require operations beyond their generation capabilities. Since our compiler constructs schedules gate by gate, the likelihood of entering such unrecoverable configurations grows with circuit depth. These results indicate that LLM-based shuttling compilation is feasible in principle but remains limited in robustness. Evaluations on previously unseen trap layouts further revealed a sharp performance degradation, indicating that the models generalize poorly beyond the architectures encountered during training.
To further improve the capabilities of our compiler, we plan to enhance the training pipeline in two stages: first using Direct Preference Optimization (DPO) [RafailovSMEMF23], and later using Gradient Regularized Policy Optimization (GRPO) [OpenAI24]. Applying DPO requires generating multiple candidate shuttling operations at each step so that the model can learn which alternatives are preferable. As discussed in Sec. 3.3, the schedule with the lowest number of operations cannot always be used as the preferred example, since following it naively may lead the model into local minima that require additional operations to escape. Instead, DPO could help the model identify invalid operations by consistently deprioritizing them. To refine this further, a lookahead strategy would be needed: for each valid candidate, the shuttlings for a certain number of next gates must be computed to determine which initial choice leads to the lowest overall number of operations. While the runtime of this approach grows exponential, it provides valuable multi-step supervision.
Building on this, GRPO offers a framework for reinforcement learning with human feedback (RLHF) [ChristianoLBMSD17]. This would require a continuous reward function that penalizes invalid schedules and rewards valid schedules with fewer operations, again ideally using a lookahead-based evaluation. Due to the large number of shuttling generations needed for DPO and the additional computational overhead of GRPO, scaling these methods will require access to more GPUs than the eight H100 GPUs available for the present work.
Overall, our work establishes the first foundation for LLM-based shuttling compilers for trapped-ion quantum computers and identifies promising directions to significantly extend their performance and reliability in future research.
@@ Acknowledgments
We acknowledge funding by the German Federal Ministry of Education and Research (BMBF) via the VDI within the project IQuAn. We thank the data center (ZDV) of the Johannes Gutenberg University for the use of their Mogon NHR and Mogon KI clusters for evaluating our LLMs.

### v2

@@ 4.2.2 Unseen architectures
In addition to linear and branched one-dimensional architectures with junctions, we evaluated the generalization capabilities of the fine-tuned LLMs on the unseen architectures shown in Fig. 5.
The layouts in 5(a) and 5(b) feature three-way junctions, similar to the branched architectures used during training. For all those architectures, none of the models generated a valid shuttling schedule for any of the evaluated circuits. While on the first architecture, all models failed immediately during schedule generation, on the second one, the models were able to execute up to four gates before failing.
In contrast, one model successfully generated a valid schedule for the four-way junction architecture shown in 5(c) for the circuit 4mod5-bdd_287. LLaMA at $T=0.7$ produced a valid schedule with 569 shuttling operations but required 55 retries, increasing the total token count by a factor of 6.0 relative to the final schedule. At $T=1.0$, the model executed 74 of the 106 gates before failing. All other models failed to generate valid schedules on this architecture. Nevertheless, partial progress was observed: Qwen-14B executed up to 52 gates, Qwen-4B up to 33 gates, and Gemma at $T=1.0$ 16 gates, while Gemma at $T=0.7$ and DeepSeek failed at an early stage.
Consistent with the results before, performance on the four-way junction architecture degraded as the number of qubits increased. For the circuit mini_alu_305, no model generated a valid schedule. LLaMA executed up to 48 gates before failing, whereas both Qwen models and Gemma executed approximately 15 gates. DeepSeek performed worst, failing to execute even a single gate. Excluding DeepSeek, these results indicate that limited progress remains possible for 10-qubit circuits, although complete schedules cannot be reliably generated.
@@
Overall, our results show that fine-tuned LLMs can generate valid shuttling schedules and can, in some cases, outperform standard heuristics. However, producing complete schedules becomes increasingly challenging with the width and depth of the algorithm. Models may enter invalid states, requiring retries that substantially inflate token counts without contributing to the final output. Generalization to previously unseen architectures remains limited. Across all experiments, LLaMA demonstrated the most stable performance, whereas Qwen, DeepSeek, and Gemma failed more frequent and at earlier stages as circuit size increased. When comparing the both Qwen models, it can be noticed that the smaller model with 4 billion parameters produced better results than the the larger model with 14 billion parameters. This indicates, together with LLaMA which is with 3 billion parameters the smallest model we tested, that smaller models can be better adapted to shuttling schedule generation than larger ones.
@@ 5 Conclusion and Outlook
We have presented a shuttling compiler for trapped-ion quantum computers based on large language models (LLMs). We fine-tuned several LLMs on datasets derived from shuttling schedules generated by classical compilers for linear and branched one-dimensional trap architectures and evaluated them on circuits of varying sizes. For small circuits on both linear and junction layouts, the models were able to generate valid schedules and in some cases even required fewer operations as compared to the standard heuristics [71, 30].
To improve the capabilities of LLM-compilers, we plan to enhance the training pipeline in two stages: first using Direct Preference Optimization (DPO) [50], and later using Gradient Regularized Policy Optimization (GRPO) [43]. Applying DPO requires generating multiple candidate shuttling operations at each step so that the model can learn which alternatives are preferable. As discussed in Sec. 3.3, the schedule with the lowest number of operations cannot always be used as the preferred example, since following it naively may lead the model into local minima that require additional operations to escape. Instead, DPO could help the model identify invalid operations by consistently deprioritizing them. To refine this further, a lookahead strategy would be needed: for each valid candidate, the shuttlings for a certain number of next gates must be computed to determine which initial choice leads to the lowest overall number of operations. While the runtime of this approach grows exponential, it provides valuable multi-step supervision.
Building on this, GRPO offers a framework for reinforcement learning with human feedback (RLHF) [8]. This would require a continuous reward function that penalizes invalid schedules and rewards valid schedules with fewer operations, again ideally using a lookahead-based evaluation. Due to the large number of shuttling generations needed for DPO and the additional computational overhead of GRPO, scaling these methods will require access to more GPUs than the eight H100 GPUs available for the present work.
Overall, our work establishes the first LLM-based shuttling compilers for trapped-ion quantum computers and may paving a way to extend performance and reliability.
@@ Acknowledgments
We acknowledge funding by the German Federal Ministry of Research, Technology and Space (BMFTR) within the project IQuAn, ATIQ and SYNQ, and the DFG SPP2514. We thank the data center (ZDV) of the Johannes Gutenberg University for the use of their Mogon NHR and Mogon KI clusters for evaluating our LLMs.